Small pieces of several metals are added to separate samples of water. Predict which metal will result in a chemical reaction. Write a balanced chemical equation for each reaction that occurs. 

  1. Copper 
  2. Calcium 
  3. Lithium 
  4. Lead
1
Expert's answer
2021-01-25T07:01:21-0500

Cu+H2O=CuO+H2

Ca+2H2O=Ca(OH)2+H2

2Li+H2O=LiOH+H2

Pb+H2O=PbO+H2


lithium will result in a chemical reaction.